17/Sep/99 Owl Shift

Hogan Nguyen (hogann@fnal.gov)
Fri, 17 Sep 1999 07:20:53 -0500 (CDT)

Shift Summary for Owl shift, 17/Sep/99

On Shift: Jianbo, Hogan

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 4

Accounting of lost time:

Runs ended this shift:


Accesses during shift:

None (Hall interlocked)

Overview of shift:

Waiting for beam

Description of Shift:

run spill time comment

1:15 KTeV Vacuum is stable at 33 uTorr (ion gaugereading)

2:00 No beam. Experts working on TeV vacuum problem.

3:55 Beam estimated to arrive in 15 minutes. Shift
crew pages BobT, Arun, Valeri, SashaL, who all
dutifully arrive.

We try to start a muon run just in case.
Initialization does not get past stage 1 due to
DOT errors on all the drives. Arun consults with

5:00 Trouble with P1 device. No estimate for beam.

6:20 MCR estimate 3 hours for beam

We can now initialize the muon run. The problem
was that the tape drives were still allocated to
onsplit. During the flush of e832 data, there was
an OCS problem which prevented onsplit from
finishing gracefully. Arun and Sydney fixed
this tape permission problem.

DC's are back at nominal HV since the chambers
have been running with good ArEth for >12 hours.

Follow up Notes on Previous Shift

Items to Follow up on Next Shift

Don't page Jason L. for frowny bananas. He'll fix it tomorrow.

Page Bob T, Sasha L, Valeri, Arun, and Ashkhan (for muon alignment) when
muons are 30 minutes away. If beam happens after noon time, contact Jason
LaDue for chamber plateau. Page Emmanuel for hyperon trigger studies
when we have hadron beam.

It would be more efficient if we just mentioned who NOT to page.